# U-net based prediction of cerebrospinal fluid distribution and ventricular reflux grading

### v1

@@ Abstract
Previous work shows evidence that cerebrospinal fluid (CSF) plays a crucial role in brain waste clearance processes, and that altered flow patterns are associated with various diseases of the central nervous system. In this study, we investigate the potential of deep learning to predict the distribution in human brain of a gadolinium-based CSF contrast agent (tracer) administered intrathecal. For this, T1-weighted magnetic resonance imaging (MRI) scans taken at multiple time points before and after intrathecal injection were utilized. We propose a U-net-based supervised learning model to predict pixel-wise signal increases at their peak after 24 hours. Its performance is evaluated based on different tracer distribution stages provided during training, including predictions from baseline scans taken before injection. Our findings indicate that using imaging data from just the first two hours post-injection for training yields tracer flow predictions comparable to those trained with additional later-stage scans. The model was further validated by comparing ventricular reflux gradings provided by neuroradiologists, and inter-rater grading among medical experts and the model showed excellent agreement. Our results demonstrate the potential of deep learning-based methods for CSF flow prediction, suggesting that fewer MRI scans could be sufficient for clinical analysis, which might significantly improve clinical efficiency, patient well-being, and lower healthcare costs.
Keywords brain clearance cerebrospinal fluid deep learning U-net glymphatic system MRI tracer
@@ 1 Introduction
Cerebrospinal fluid (CSF) and its elementary roles in cushioning the brain and spinal cord, facilitating buoyancy to prevent nerve roots or blood vessels being compressed and transporting nutrients, have been known for a long time. Recent research findings however dramatically expanded our understanding of CSF, particularly its critical function in brain solute clearance and biomarker transportation(Fame and Lehtinen, 2020; Wichmann et al., 2022; Thomas, 2019; Bjerke and Engelborghs, 2018; Bohr et al., 2022). The role of CSF in clearing proteins has gained considerable attention, given the main characteristic of neurodegenerative diseases to be abnormal protein aggregation, such as amyloid- or tau in Alzheimer’s disease (AD) and -synuclein in Parkinson’s disease (Pingle et al., 2023; Nedergaard and Goldman, 2020). Apart from that, CSF plays a crucial role in the diagnosis, monitoring, and understanding of autoimmune, infectious or neuropsychiatric disorders as well as hydrocephalus, idiopathic intracranial hypertension, ischemic stroke, and brain cancers, indicating that changes in fluid dynamics can influence tumor progression and treatment outcomes (Jin and Munson, 2023; Honnorat and Plazat, 2018). Furthermore, recent scientific findings suggest that CSF helps regulate the sleep-wake cycle, with sleep deprivation linked to disturbed flow dynamics (Eide et al., 2021a; Vinje et al., 2023). Distorted clearance mechanisms may contribute to the pathogenesis and progression of these conditions, thus an enhanced understanding of CSF dynamics is expected to offer valuable insights into the mechanisms by which molecules are exchanged between the blood, brain, and cerebrospinal fluid, and may ultimately open new therapeutic avenues (Eide et al., 2021b). Among the various imaging techniques employed for brain clearance analysis, intrathecal contrast-enhanced (CE) MRI methods haven been well-established in clinical practice(Ringstad et al., 2017). They permit in vivo tracking of an exogeneous tracer to mirror the behavior of endogenous solutes within the brain, but necessitate a series of MRI scans, which results in inpatient admission, and consequently, significant healthcare costs. Previous research on CSF analysis has used various mathematical and computational models to understand its dynamics and relationship with neurological conditions. However, these models often depend on advanced computational resources or medical assumptions made in the modeling process(Valnes et al., 2020; Zapf et al., 2022; Hornkjøl et al., 2022; Zeinhofer and Mardal, 2024). Given that distribution patterns vary considerably across different neurodegenerative conditions(Ringstad et al., 2017; Eide et al., 2020) and the increasing availability of data, a more robust and flexible approach should be employed. In that context, data driven machine learning, and particularly deep learning could be a promising alternative incorporating hidden behaviors observed only in data. Despite the advancements in both traditional and machine learning techniques, there remains a gap in the direct application of deep learning to forecast CSF flow. In this study, we investigate the use of a convolutional neural network (CNN) in tracer analysis and evaluate how well deep learning-based methods are able to predict CSF tracer distribution patterns. Besides evaluating image similarity of predicted MRI scans compared to real imaging data using evaluation metrics common in deep learning, we additionally use a clinically oriented reflux assessment scheme as a method to judge the model’s performance. This is relevant because assessment of ventricular reflux has been found clinically useful in the diagnostic pre-operative work-up of idiopathic normal pressure hydrocephalus (iNPH) (Eide et al., 2020, 2021b). The present findings demonstrate the use of deep learning techniques in CSF distribution prediction and indicate that repeated MRI may be avoided, resulting in reduced medical cost, consumed clinical resources, as well as patient burden.
@@ 2 Related work
Previous research on cerebrospinal fluid analysis has employed various mathematical and computational modeling approaches to understand its dynamics and interplay with several neurological conditions. CNNs, for example, were utilized numerous times to predict AD from brain MRI data, though almost exclusively not contrast-enhanced. Pathological molecular hallmarks of AD, such as amyloid--42 and tau proteins, were quantified using an ensemble-based model(Popuri et al., 2020). Schweizer et al. examined the potential to use CNNs for classification of individual cells that CSF contains(Schweizer et al., 2023), and real-time CSF flow velocities were estimated using inflow effect-mediated signal increase of in vivo functional MRI(Diorio et al., 2024). Four-dimensional flow of CSF has been assessed with various patient cohorts, fluid markers, and regions of interest(Rivera-Rivera et al., 2024), although the limitations of this imaging approach due to low velocities associated with CSF motion and large T1 relaxation values were continuously faced. While well-established mathematical approaches such as mesh-based finite element methods and partial differential equation-constrained optimization were used to directly model diffusion of CSF in the human brain(Valnes et al., 2020), they face some drawbacks such as high computational cost and the need of high-quality meshes to simulate complex geometries. Quite recently, physics-informed neural networks (PINNs) were introduced to integrate physical laws into neural networks for various applications, including heat transfer and diffusion. Zapf et al.(Vinje et al., 2023; Zapf et al., 2022) highlighted the potential of PINNs in solving inverse problems governed by partial differential constrained optimization, which were extended to estimate the apparent diffusion coefficient of cerebrospinal fluid tracers in specific regions of interest in the brain (subcortical white matter subregiona and full brain parenchyma) and experimentally validated with MRI data from sleep-deprived patients. Their findings suggest that diffusion alone may not be sufficient to model molecule distribution or clearance in the brain, supporting the glymphatic model’s perspective on molecular transport. In the domain of computational fluid dynamics, physiological boundary conditions improved the simulations of intracranial pressures and flow velocities through the cerebral aqueduct and spinal subarachnoid space (SAS) (Vandenbulcke et al., 2022). While the aforementioned approaches each offer insights into cerebrospinal fluid modeling, they all rely on specific assumptions and simplifications that may not universally represent the complex process of CSF distribution and clearance across different conditions; thus, in our study, we employ deep learning techniques, which do not require prior information about brain anatomy, CSF production, circulation, or volume properties, aiming to offer a more generalized solution.
@@ 3.1 Study population
This was a single-center retrospective study that utilized MRI data from a cohort of 136 patients of different neurological conditions. Table 1 shows the distribution of neuropathological findings for all of the analyzed data.
| N | Age | Sex (Female/Male) | Body mass index | |
@@ 3.2 Injection protocol, MRI hardware, and data preprocessing
To evaluate the enrichment of a CSF tracer in the human brain, the MRI contrast agent gadobutrol was utilized as CSF tracer. An interventional neuroradiologist administered an intrathecal injection of gadobutrol at a dosage of ( of gadobutrol; Gadovist, Bayer Pharma AG, Germany). T1-weighted MRI scans were performed before and at various intervals post-injection. For practical reasons, MRI examinations could not be conducted at the exact same times for each study participant, but there is consistency grouping the scans into the following time intervals: 0 hours (before injection, baseline), 1-2 hours, 3–5 hours, 5-6 hours, 6-10 hours, 24 hours, 48 hours, and 29 days (full clearance, control scan). Some patients were scanned multiple times during an interval. We used a 3 Tesla Philips Ingenia MRI scanner (Philips Medical Systems) with equal imaging protocol settings at all time points to acquire sagittal 3D T1-weighted scans. The imaging parameters were: repetition time = ‘shortest’ (typically 5.1 ms), echo time = ‘shortest’ (typically 2.3 ms), flip angle=8 degrees, field of view and matrix = pixels (reconstructed ). The slices were sampled 184 overcontiguous (overlapping) with 1 mm thickness that were automatically reconstructed to 368 slices with 0.5 mm thickness; total duration of each image acquisition was 6 min and 29 s. To secure consistency and reproducibility of the MRI slice placement and orientation, slice orientation of image stacks was defined using an automated anatomy recognition protocol based on landmark detection in MRI data (SmartExamTM, Philips Medical Systems) for every time point. The images for each patient was registered/aligned using FreeSurfer software (version 6.0) (http://surfer.nmr.mgh.harvard.edu/) and re-sampled to the FreeSurfer standard of voxels with voxel size of (Eide et al., 2021a, b; Ringstad et al., 2018). The images was normalized by dividing the signal values with the mean signal value of a reference volume within the intraconal orbital fat (Eide et al., 2021a). Then, each images were centered and the sagittal and axial slices were exported as matrices. For predicting tracer distribution after 24 hours, we fed our model with different information during training and compared the results. For grayscale images, possible intensities range from 0 (black) to 1 (white) and in between different shades of gray, as shown in Figure 1 for illustrative purposes.
3.3 Distribution and clearance pattern
@@
A series of T1-weighted images of the patient taken after intrathecal tracer administration, exemplified in Fig. 2, shows the tracer transport and elimination mechanisms. Following the injection of gadobutrol into the spinal canal at the lumbar region, the tracer ascends through the spinal canal, and typically reaches the subarachnoid space (SAS) at the level of the foramen magnum (Liu et al., 2022; Dreha-Kulaczewski et al., 2015). Significant spinal resorption means that only about of the injected tracer reaches intracranial space (Vinje et al., 2023; Eide et al., 2021c). Once in the cisterna magma, the tracer rapidly moves to subarachnoid cisterns and then distributes along major arteries within the cerebral fissures. To evaluate the enrichment of a CSF tracer in the human brain, the MRI contrast agent gadobutrol was utilized as CSF tracer, showing significant accumulation in areas adjacent to the major arteries (Vinje et al., 2023; Ringstad et al., 2018). MRI scans performed 24 hours after injection reveal diffuse tracer enhancement around the brain, with signal enhancement peaking typically at 24 hours after injection (Ringstad et al., 2017). Gadobutrol clearance is driven by resorption initially from the spinal canal, and peak gadobutrol concentration in the blood occurs around five to six hours after injection (Eide et al., 2021c; Hovd et al., 2022; Shah et al., 2023; Melin et al., 2023). At the four-week post-injection mark, MRI scans reveal no evidence of contrast agent accumulation in the brain. Still, the exact underlying mechanisms of tracer distribution and clearance remain not fully understood (Plog and Nedergaard, 2018; Iliff et al., 2012).
@@ 3.4 Model architecture
where is a given loss function and is the training data. There exists a variety of loss functions. Here, we use a regression model and compare the two commonly chosen loss functions: the -loss defined as , and the -loss, defined as . Experimenting with different loss functions is a crucial part of designing a deep learning model, because the choice of the loss function significantly impacts the model’s performance, behavior, and convergence during training, which has been shown a number of times (Shan et al., 2023). Figure 3 depicts a simplified illustration of the training process.
@@
In this study, we used a U-net architecture for the prediction of the tracer distribution. Introduced by Olaf Ronneberger et al. in 2015 (Ronneberger et al., 2015), U-net is a model originally proposed for 2D or 3D image segmentation, particularly in biomedical applications (Cai and Yuan, 2022; Siddique et al., 2021). Thereafter, it served as a blueprint for a plethora of variations that have so far been proposed and used in several application areas (Wang et al., 2023; Azad et al., 2022; Pai and M., 2023), which all share a common structure inspired by an encoder-decoder architecture comprising two primary paths: the contracting (or compressive) path and the expansive (or up-sampling) path. The contracting path functions similarly to a convolutional neural network (CNN), where images are down-sampled at first using a convolutional kernel followed by an activation function. Conversely, the expansive path up-samples the compressed representation to reconstruct the image. In addition, skip connections link the corresponding layers of the contracting and expansive paths, helping to preserve spatial information that is typically lost during the down-sampling process (Wilm et al., 2024). In contrast to other supervised deep learning methods such as CNNs and fully convolutional neural networks (FCNs), U-nets have a more complicated architecture, yet have been shown to perform better, particularly on medical imaging data and small data sets (Wang et al., 2023; Azad et al., 2022). Furthermore, U-Net-based models are well-equipped to handle small structures and objects within an image, a capability crucial for accurately predicting tracer fluid distribution. For the implementation of our 2D U-Net architecture, we used PyTorch, a variable number of input channels, 64 features, 8 batch size, Adam optimizer, 200 or 250 epochs, learning rate, 3×3 convolutional kernels followed by batch normalization and rectified linear unit as activation function, and 2×2 maximal pooling size. Batch normalization mitigates the internal covariate shift by normalizing layer inputs, stabilizing and accelerating the training process. It allows higher learning rates, improves the gradient flow, and has a regularizing effect. ReLU, defined as
@@
introduces a non-linearity into the network, which is crucial for modeling complex, non-linear relationships. ReLU offers a computationally efficient non-linear transformation, and enhances the network’s efficiency by activating only a subset of neurons, which can lead to a reduced computational load and mitigate the risk of overfitting. Overfitting occurs when a model becomes too tailored to the training data, capturing noise and specific patterns to the extent that it negatively impacts its performance on new, unseen data. Additionally, ReLU helps alleviate the vanishing gradient problem, particularly in deeper networks, by ensuring that gradients remain significant, thereby facilitating more robust and efficient training. Firstly, we use a U-net as illustrated in Figure 4(a) with two input channels (for sagittal and axial scans, respectively) to use scans from baseline (before injection) as training input. With this, we address how well our model may accurately predict tracer distribution after 24 hours providing it with specific patient anatomy only. Changing the number of input channels allows the model to further use tracer-enriched scans. By comparing the model performance of varying training data, we aim to assess to what extent different tracer distribution stages contribute to better prediction performance. In any model training, the available data is split into cases for training and the remaining for testing. Figure 4(b) provides an overview of the data processing workflow, deep-learning model usage, and clinical interpretation.

### v2

@@ Abstract
Previous work indicates evidence that cerebrospinal fluid (CSF) plays a crucial role in brain waste clearance processes, and that altered flow patterns are associated with various diseases of the central nervous system. In this study, we investigate the potential of deep learning to predict the distribution in human brain of a gadolinium-based CSF contrast agent (tracer) administered intrathecal. For this, T1-weighted magnetic resonance imaging (MRI) scans taken at multiple time points before and after injection were utilized. We propose a U-net-based supervised learning model to predict pixel-wise signal increase at its peak after 24 hours. Performance is evaluated based on different tracer distribution stages provided during training, including predictions from baseline scans taken before injection. Our findings show that training with imaging data from only the first two hours post-injection yields tracer flow predictions comparable to models trained with additional later-stage scans. Validation against ventricular reflux gradings from neuroradiologists confirmed alignment with expert evaluations. These results demonstrate that deep learning-based methods for CSF flow prediction deserve more attention, as minimizing MR imaging without compromising clinical analysis could enhance efficiency, improve patient well-being, and lower healthcare costs.
@@ 1 Introduction
The cerebrospinal fluid (CSF) flow and its role in solute transport and clearance has received a lot of attention since the introduction of the glymphatic system in 2012 [1, 2]. In particular, the CSF and its interaction with the extra-cellular environment of the brain through perivascular spaces (PVSs) may be crucial for the clearance of metabolic waste, which is known to accumulate in Alzheimer’s and Parkinson’s diseases [3], to name a few. However, the CSF dynamics is complex as it is dictated by not only cardiac pulsations, but respiration [4, 5], vasomotion [6, 7], sleep waves [8, 9] as well as the circadian rhythm [10]. Hence, the governing mechanisms are only partially known, and modeling attempts have as such only partially explained the tracer distribution in humans [11, 12]. Furthermore, neuroanatomic characteristics can vary significantly between individuals, and differences in tracer distributions may be driven mostly by geometry difference. Given the partially unknown physical processes, it is interesting to investigate whether purely data-driven approaches can predict future tracer distributions and what types of data yield the best predictions.
Several novel imaging methods have been proposed for evaluating glymphatic function, including DTI-ALPS [13] and imaging of enlarged PVSs [14], which focus on structural features. Furthermore, flow imaging for cardiac or respiratory pulsations is available with phase-contrast MR [15, 16], both in 2D slices and in 3D [17]. These techniques, in particular, have been shown to be useful for the ventricular system and the cranio-cervical junction in conditions such as normal pressure hydrocephalus (NPH) [18, 19] and the Chiari malformation [20], where the abnormal CSF pulsations at cardiac frequencies are used in diagnostics. However, it is not clear that such imaging is useful for solute transport and clearance that may be governed by slower processes. As such, one new and promising method is exploiting intrathecal MR contrast to visualize the slow transport of gadobutrol in the CSF and the extracellular space of the brain [21, 22]. A main advantage of the technique is that imaging illustrates the solute transport on the long time-scale of days and it has been demonstrated that there is great variation among patients [23], and that the transport is reduced in sleep-deprived individuals [24]. A major drawback of the procedure is that it is time-consuming, as patients undergo multiple imaging sessions over several days. Furthermore, only a few centers worldwide currently perform intrathecal MR contrast investigations, as the methodology is off-label (although it is occasionally used for assessing CSF leaks in spontaneous intracranial hypotension [25]).
Modeling and/or data-driven approaches could significantly reduce patient scanning time if they provide accurate predictions. Furthermore, since the procedure is off-label, one may question whether tracer distributions can be learned solely from anatomy, as seen in a standard T1-weighted image.
In this study, we adapt a modification of U-net [26], a specific type of convolutional neural network (CNN), to tracer analysis and hereby evaluate the ability of deep learning methods to predict CSF distribution patterns of 2D MR imaging data. In addition to evaluating the image similarity of predicted scans compared to real imaging data using evaluation metrics common in deep learning, we additionally use a clinically oriented reflux assessment scheme as a method to judge the model performance. This is significant because ventricular reflux assessment has been clinically useful in the preoperative diagnostic workup of idiopathic normal pressure hydrocephalus [27, 23]. The present findings showcase that even basic machine-learning techniques can be used to model complex physiological processes, and more specifically and uniquely in this context, demonstrate the use of deep learning techniques in CSF distribution prediction. In the future, it is hoped that this approach could avoid the need for repeated MRIs, thereby reducing medical costs, clinical resource consumption, and patient burden.
@@ 2 Related work
Previous research on cerebrospinal fluid analysis has employed various mathematical and computational modeling approaches to understand its dynamics and interplay with several neurological conditions. CNNs, for example, were utilized numerous times to predict AD from brain MRI data, though almost exclusively not contrast-enhanced. Pathological molecular hallmarks of AD, such as amyloid-β-42 and tau proteins, were quantified using an ensemble-based model[28]. Schweizer et al. examined the potential to use CNNs for classification of individual cells that CSF contains[29], and real-time CSF flow velocities were estimated using inflow effect-mediated signal increase of in vivo functional MRI[30]. Four-dimensional flow of CSF has been assessed with various patient cohorts, fluid markers, and regions of interest[31], although the limitations of this imaging approach due to low velocities associated with CSF motion and large T1 relaxation values were continuously faced.
While well-established mathematical approaches such as mesh-based finite element methods and partial differential equation-constrained optimization were used to directly model diffusion of CSF in the human brain[11], they face some drawbacks such as high computational cost and the need of high-quality meshes to simulate complex geometries. Quite recently, physics-informed neural networks (PINNs) were introduced to integrate physical laws into neural networks for various applications, including heat transfer and diffusion. Zapf et al.[32] highlighted the potential of PINNs in solving inverse problems governed by partial differential constrained optimization, which were extended to estimate the apparent diffusion coefficient of cerebrospinal fluid tracers in specific regions of interest in the brain (subcortical white matter subregions and full brain parenchyma) and experimentally validated with MRI data from sleep-deprived patients[12]. Their findings suggest that diffusion alone may not be sufficient to model molecule distribution in the brain, supporting the glymphatic model’s perspective on molecular transport. In the domain of computational fluid dynamics, physiological boundary conditions improved the simulations of intracranial pressures and flow velocities through the cerebral aqueduct and spinal SAS[33].
In this study, we explore the application of deep learning techniques to predict CSF tracer distribution; an approach that has not been previously applied to this problem. By doing so, we aim to demonstrate the potential of a simple data-driven machine-learning framework for modeling complex physiological processes, without explicitly requiring anatomical modeling of CSF production, circulation, or volume properties.
@@ 3.1 Study population
This work was a single-center retrospective study that utilized MRI data from a cohort of 136 patients with different neurological conditions. All participants provided informed consent prior to their involvement in the study. Table 1 shows the distribution of neuropathological findings for all of the analyzed data. This distribution of conditions reflects the typical case mix observed in our clinical practice for individuals requiring CSF assessment, underscoring its representativeness for this specific patient population at our center. For training and testing our model, we randomly split the Ntotal = 136 cases in Ntrain = 105 and Ntest = 31.
| Ntotal | Age | Sex (Female/Male) | Body mass index | |
@@ 3.2 Injection protocol, MRI hardware, and data preprocessing
To evaluate the enrichment of a CSF tracer in the human brain, the MRI contrast agent gadobutrol was utilized. An interventional neuroradiologist administered an intrathecal injection of gadobutrol at a dosage of ( of gadobutrol; Gadovist, Bayer Pharma AG, Germany). T1-weighted MRI scans were performed before and at various intervals post-injection. For practical reasons, MRI examinations could not be conducted at the exact same times for each study participant, but there is consistency grouping the scans into the following time intervals: 0 hours (before injection, baseline), 1-2 hours, 3–5 hours, 5-6 hours, 6-10 hours, 24 hours, 48 hours, and 29 days (full clearance, control scan).
@@
We used a 3 Tesla Philips Ingenia MRI scanner (Philips Medical Systems) with equal imaging protocol settings at all time points to acquire sagittal 3D T1-weighted scans. The imaging parameters were: repetition time = ‘shortest’ (typically ), echo time = ‘shortest’ (typically ), flip angle=, × field of view and matrix of size 256×256 (reconstructed 512×512). The slices were reconstructed with thickness that were automatically reconstructed to 368 slices with distance; total duration of each image acquisition was and . To secure consistency and reproducibility of the MRI slice placement and orientation, slice orientation of image stacks was defined using an automated anatomy recognition protocol based on landmark detection in MRI data (SmartExamTM, Philips Medical Systems) for every time point. The images for each patient were registered/aligned using FreeSurfer software (version 6.0) (http://surfer.nmr.mgh.harvard.edu/) and re-sampled to the FreeSurfer standard of voxels with voxel size of ××[24, 23, 34]. The images was normalized by dividing the signal values with the mean signal value of a reference volume within the intraconal orbital fat[24]. Then, images were each centered and the sagittal and axial slices were exported as 256×256 matrices. In this study, intensity scaling between 0 and 1 was performed independently for each 2D slice to normalize intensity variations within slices, ensuring consistent signal distribution for model training and analysis; however, as this approach may inadvertently introduce artifacts in non-enhanced tissue intensities, and alternative global normalization strategies will be explored in future work. For predicting tracer distribution after 24 hours, we fed our model with different information during training and compared the results.
Importantly, for each patient and time point, we used one 2D scan slice in the axial plane and one in the sagittal plane as in-/output to the network, which we refer to as an image pair in the following. Predicted sagittal and transverse slices provide complementary views of the ventricular system, and both planes are valuable for grading ventricular reflux. The sagittal slice is particularly well-suited for assessing CSF tracer enhancement in the third ventricle relative to the suprasellar cistern below, determining whether the enhancement is equal or differs. In contrast, the axial slice is ideal for evaluating enhancement within the ventricles and the corresponding level of the subarachnoid space outside the cerebral convexities. In principle and given sufficient computational resources, our approach can be extended to volumetric data.
3.3 Distribution pattern
@@
A series of T1-weighted images of the patient taken after intrathecal tracer administration, exemplified in Fig. 1, show the tracer transport and elimination mechanisms. Following the injection of gadobutrol into the spinal canal at the lumbar region, the tracer ascends through the spinal canal, and typically reaches the subarachnoid space (SAS) at the level of the foramen magnum[35]. Significant spinal resorption means that only about of the injected tracer reaches intracranial space[12]. Once in the cisterna magma, the tracer rapidly moves to subarachnoid cisterns and then distributes along major arteries within the cerebral fissures. To evaluate the enrichment of a CSF tracer in the human brain, the MRI contrast agent gadobutrol was utilized as CSF tracer, showing significant accumulation in areas adjacent to the major arteries[12]. MRI scans performed 24 hours after injection reveal diffuse tracer enhancement around the brain, with signal enhancement peaking typically at 24 hours after injection[22]. Gadobutrol clearance is driven by resorption initially from the spinal canal, and peak gadobutrol concentration in the blood occurs around five to six hours after injection[36, 37, 38, 39]. At the four-week post-injection mark, MRI scans reveal no evidence of contrast agent accumulation in the brain. Still, the exact underlying mechanisms of tracer distribution and clearance remain not fully understood [40].
@@ 3.4 Model architecture
where is a given loss function and is the training data. There exists a variety of loss functions. Here, we use a regression model and compare the two commonly chosen loss functions: the -loss defined as , and the -loss, defined as . Experimenting with different loss functions is a crucial part of designing a deep learning model, because the choice of the loss function significantly impacts the model’s performance, behavior, and convergence during training, which has been shown a number of times [41]. Figure 2 depicts a simplified illustration of the training process.
@@
In this study, we used a U-net architecture for the prediction of the tracer distribution. Introduced by Olaf Ronneberger et al. in 2015[26], U-net is a model originally proposed for 2D or 3D image segmentation, particularly in biomedical applications[42, 43]. Thereafter, it served as a blueprint for a plethora of variations that have so far been proposed and used in several application areas[44, 45], which all share a common structure inspired by an encoder-decoder architecture comprising two primary paths: the contracting (or compressive) path and the expansive (or up-sampling) path. The contracting path functions similarly to a convolutional neural network (CNN), where images are down-sampled at first using a convolutional kernel followed by an activation function. Conversely, the expansive path up-samples the compressed representation to reconstruct the image. In addition, skip connections link the corresponding layers of the contracting and expansive paths, helping to preserve spatial information that is typically lost during the down-sampling process[46]. In contrast to other supervised deep learning methods such as CNNs and fully convolutional neural networks (FCNs), U-nets have a more complicated architecture, yet have been shown to perform better, particularly on medical imaging data and small data sets[44]. Furthermore, U-Net-based models are well-equipped to handle small structures and objects within an image, a capability crucial for accurately predicting tracer fluid distribution.
@@
For the implementation of our 2D U-Net architecture, we used PyTorch, a variable number of input channels, 64 features, 8 batch size, Adam optimizer, 200 or 250 epochs, learning rate, 3×3 convolutional kernels followed by batch normalization and rectified linear unit as activation function, and 2×2 maximal pooling size. Firstly, we use a U-net as illustrated in Figure 3(a) with two input channels (for sagittal and axial scans, respectively) to use scans from baseline (before injection) as training input. With this, we address how well our model may accurately predict tracer distribution after 24 hours providing it with specific patient anatomy only. Changing the number of input channels allows the model to further use tracer-enriched scans. By comparing the model performance of varying training data, we aim to assess to what extent different tracer distribution stages contribute to better prediction performance. Figure 3(b) provides an overview of the data processing workflow, deep-learning model usage, and clinical interpretation.